OpenStudy (anonymous):

What is the best first step to solve the system using elimination? x - 5y = 4 3x + 7y = -17

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Nope, I have no idea how to do this

OpenStudy (skullpatrol):

What is the plan of attack with using "elimination"?

OpenStudy (anonymous):

Identify and create opposite coefficients

OpenStudy (jdoe0001):

yes

OpenStudy (skullpatrol):

Yep, so look for opposites first.

OpenStudy (anonymous):

multiply fist equation by 3 and then substract them

OpenStudy (skullpatrol):

If you can't find any multiply both sides of the equation to "create" them.

OpenStudy (anonymous):

-3(x-5y=4) *multiply the top equation by -3 so the x values will be opposites 3x+7y=-17 -3x+15y=-12 3x+ 7y=-17 *then simply add the top equation to the bottom one 22y=-29 y=\[-\frac{ 29 }{ 22 }\]